Hello
Please solve . Given the functions f(x) =5x + 17 and g(x)= 6x^2 -7x-3 find (f/g)(x) and the domain of f/g.
Sorry about the mixup.
Thanking you

Answer by stanbon(75887)

Given the functions f(x) =5x + 17 and g(x)= 6x^2 -7x-3 find (f/g)(x) and the domain of f/g.
------------
f/g = f(x)/g(x) = (5x+17)/[(2x-3)(3x+1)]
Domain of f/g is "all Real Numbers except x = 3/2 and x = -1/3
======================
Cheers,
Stan H.